OpenStudy (anonymous):

write using exponent. -3 . m . 4 . n . n

OpenStudy (anonymous):

-3m.(2n)^2

Parth (parthkohli):

\( \color{Black}{\Rightarrow -3 \times m \times -4 \times n \times n }\) \( \color{Black}{\Rightarrow -12 \times m \times n^2}\) \( \color{Black}{\Rightarrow -12mn^2 }\)

Parth (parthkohli):

Or: \( \color{Black}{\Rightarrow -3 \times m \times 2^2 \times n^2}\) \( \color{Black}{\Rightarrow -3m \times (2n)^2}\)

Parth (parthkohli):

There are tons of ways to do this - which one do you ask>